OpenStudy (anonymous):

Solve the system, using substitution. Medal for correct steps, and for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

I don't want to type out in words what i am doing, but try to follow along with the math. a + b = 4 2a + 3b = 11 a = -b + 4 2(-b + 4) + 3b = 11 -2b + 8 + 3b = 11 b = 3 a + (3) = 4 a = 1 hope it helped

OpenStudy (anonymous):

forgot pairs (1,3)

hero (hero):

KaupKel, make sure you double check that previous question to ensure you have the correct answer.

OpenStudy (anonymous):

Ok.